A 30-cm-diameter, 10-cm-high vertical cylindrical vessel equipped with a vertical tube at the edge is rotated about its vertical axis at a constant angular velocity of 15 rad/s. If the water rise in the tube is 30 cm, determine the net vertical pressure force acting on the vessel.

1
Expert's answer
2021-11-03T13:00:25-0400



net vertical pressure force acting on the vessel:


"\Delta P=\rho gh-\rho \omega^2R^2/2=1000(9.8\cdot0.3-15^2\cdot 0.15^2/2)=409.75" Pa